# Improving Retrieval in Sponsored Search by Leveraging Query Context Signals

### v1

@@ 2.2 Model Architecture & Training
$\mathcal{L}^{D}(\theta,\mathcal{B})$ $=\sum_{Q,K\in\mathcal{B}}\log(\frac{\exp(\mbox{Sim}(Q,K))}{\sum_{K^{\prime}\in\mathcal{B}}\exp(\mbox{Sim}(Q,K^{\prime}))})$
@@ 3.5 Online A/B Testing
To validate the effectiveness of Augmented Unity in a real-world setting, we conducted extensive online A/B testing on live traffic of a prominent commercial search engine, spanning over 160 countries. We deployed both the NLG and DR components of Augmented Unity and compared their performance against an ensemble of state-of-the-art retrieval techniques, including proprietary DR and NLG models, Unity, large language models, extreme classification, and graph-based methods. We measured the overall revenue, ad clicks, Quick Back Rate, and ad defect. Quick Back Rate (QBR) denotes the percentage of ad clicks with users quickly returning to the search results page. Ad defect, measured by offline relevance models, denotes the percentage of irrelevant ads shown to users. Table 7 summarizes the online A/B testing results, segmented by English and non-English queries. Augmented Unity improved user engagement, yielding statistically significant increases in overall clicks – a 0.72% lift for non-English queries and a 0.33% lift for English queries. Critically, these gains were not accompanied by any statistically significant degradation in QBR or ad defect. This suggests that Augmented Unity was able to retrieve keywords that aligned with the user intent. As a result of the improved user engagement, we observed substantial revenue gains – 1.43% for non-English queries and 1.02% for English queries – underscoring the tangible business impact of our approach within a real-world sponsored search ecosystem.

### v2

@@ 2.1 Query Context Signals
We utilize various query normalization techniques, including spell correction, to map minor query variations to a canonical form for cache storage and lookups. With these normalizations, our cache hit rate is approximately 70% of all user search requests.
@@ 2.2 Model Architecture & Training
$\mathcal{L}^{D}(\theta,\mathcal{B})$ $=\sum_{Q,K\in\mathcal{B}}\log\frac{\exp(\mbox{Sim}(Q,K))}{\sum_{K^{\prime}\in\mathcal{B}}\exp(\mbox{Sim}(Q,K^{\prime}))}$
@@ 3.5 Online A/B Testing
Decile $\Delta$ Query Coverage $\Delta$ Ad Impressions
English Non-English English Non-English
1 0.66% 1.49% 0.84% 1.71%
2 0.09% 0.70% 0.01% 0.35%
3 1.16% 0.48% 0.37% 0.86%
4 0.49% 0.51% 0.26% 0.48%
5 0.41% 0.51% 0.59% 0.61%
6 0.35% 0.56% 0.55% 0.70%
7 0.61% 0.71% 0.66% 0.54%
8 0.76% 0.69% 0.98% 0.74%
9 0.51% 1.12% 1.47% 1.09%
10 1.07% 0.76% 1.66% 0.76%
Table 8: Percentage Change in query coverage and Ad impression for different deciles in online A/B tests on sponsored search
To validate the effectiveness of Augmented Unity in a real-world setting, we conducted extensive online A/B testing for 30 days on live traffic of a prominent commercial search engine, spanning over 160 countries. We deployed both the NLG and DR components of Augmented Unity and compared their performance against an ensemble of state-of-the-art retrieval techniques, including proprietary DR and NLG models, Unity, large language models, extreme classification, and graph-based methods. We measured the overall revenue, ad clicks, Quick Back Rate, and ad defect. Quick Back Rate (QBR) denotes the percentage of ad clicks with users quickly returning to the search results page. Ad defect, measured by offline relevance models, denotes the percentage of irrelevant ads shown to users. Table 7 summarizes the online A/B testing results, segmented by English and non-English queries. Augmented Unity improved user engagement, yielding statistically significant increases in overall clicks – a 0.72% lift for non-English queries and a 0.33% lift for English queries. Critically, these gains were not accompanied by any statistically significant degradation in QBR or ad defect. This suggests that Augmented Unity was able to retrieve keywords that aligned with the user intent.
We also analyzed our online A/B experiment results by grouping queries into frequency-based deciles. Decile 1 contains highly frequent queries, while decile 10 consists of a large number of rare queries. Table 8 shows the query coverage (the fraction of queries for which any sponsored content was shown) and ad impressions (the total number of ads displayed). We observe an increase in both query coverage and ad impressions across all deciles, with particularly strong gains on tail queries, which are often longer and more ambiguous. As a result of the improved user engagement, we observed substantial revenue gains – 1.43% for non-English queries and 1.02% for English queries – underscoring the tangible business impact of our approach within a real-world sponsored search ecosystem.
@@ Appendix B Multilingual Analysis
Appendix C Error Analysis
Example: Noise from Web Snippets
Query compensation injury claim
Web Snippet Report an occupational injury to the Compensation Fund Report injuries sustained by your employees in the course of their work to the Compensation Fund within seven days of the occurrence. The fund covers permanent, casual workers, trainees and apprentices who are injured in the course of their work and lose income or are impaired as a result
Predictions
Augmented Unity what to claim on income tax
Unity compensation for injury
Analysis
The model incorrectly picked up the token “income” from the context, leading to an irrelevant keyword related to income tax.
Example: Location Bias
Query cruises singles
Web Title Singles Cruise Deals | Marella Cruises | TUI.co.uk
Predictions
Augmented Unity cruises holidays uk
Unity cruises for singles only
Analysis
The cached web result was location-specific (UK), leading to a geographically biased keyword retrieval.
Table 10: Some qualitative examples of failure modes for Augmented Unity compared to Unity.
While Augmented Unity leads to significantly better model performance overall, there are a few cases where adding context can lead to errors. Table 10 shows a few selected failure modes for the model.
Appendix D Alternatives to GPT-4 for Query Profile
Model Precision@10
Avg Novelty
(1-5)
Avg Diversity
(1-5)
GPT-4 6.9 2.19 2.53
Finetuned Mistral-7b 7.62 2.05 2.36
Table 11: Performance of Finetuned Mistral-7b vs GPT-4 for QProfile Rewrites
Running GPT-4 for a large number of queries to build a sufficiently large Query Profile cache can get fairly expensive. In order to mitigate this, we experiment with a finetuning based approach with Mistral 7B Jiang et al. (2023) with LoRA Hu et al. (2022) for generating QProfile Rewrites. We evaluate the generated rewrites on three metrics precision, novelty and diversity. Precision is calculated as described in 3.3, while novelty and diversity are calculated by GPT-4, where we get a score between 1-5 for novelty for each of rewrites and a score between 1-5 for diversity for all 10 rewrites together. Table 11 shows a comparison of results between rewrites generated by GPT-4 and Mistral-7b. Our results show that fine-tuned Mistral-7B achieves comparable performance to GPT-4 in terms of query rewrite precision, at the cost of novelty and diversity, while significantly reducing inference costs. Improving novelty and diversity without losing out on precision could be a fruitful direction for a future work.